Leighann15:

Que pressure may be represent as a function of height above the surface of the Earth as a shown below P(h)=Poe^-.00012h

Vocaloid:

I believe this is the full question? Air pressure may be represented as a function of height above the surface of the Earth as shown below. P(h)=PoE^(-.00012h) In this function, P0 is air pressure at sea level, and h is measured in meters. Which of the following equations will find the height at which air pressure is 65% of the air pressure at sea level? A. Po=.65PoE^-.00012h B. .65Po=PoE^-.00012h C. .65=h*e^-.00012 D. h=.65e^-.00012

Vocaloid:

P0 is the pressure at sea level P(h) is the pressure at some height h so if the pressure is equal to 65% of the air pressure at sea level, naturally, P(h) = 0.65*P0. Substitute 0.65*P0 for P(h) and you should have your final equation.